Write a sine function with the given amplitude, period, and vertical shift.
amplitude: 4
period: 2/3pi
vertical shift: -2

Answer by lwsshak3(11628)

Write a sine function with the given amplitude, period, and vertical shift.
amplitude: 4
period: 2/3pi
vertical shift: -2
***
Form of equation for sin function: y=Asin(Bx-C)+D, Amplitude=A, Period=2π/B, Phase shift=C/B, D=Vertical shift
For given problem:
Amplitude,A=4
Period=2π/3=2π/B
B=3
C=0
D=-2
Equation: y=4sin(3x)-2